# World-Env: Leveraging World Model as a Virtual Environment for VLA Post-Training

### v3

@@ Abstract
Vision-Language-Action (VLA) models trained via imitation learning suffer from significant performance degradation in data-scarce scenarios due to their reliance on large-scale demonstration datasets. Although reinforcement learning (RL)-based post-training has proven effective in addressing data scarcity, its application to VLA models is hindered by the non-resettable nature of real-world environments. This limitation is particularly critical in high-risk domains such as industrial automation, where interactions often induce state changes that are costly or infeasible to revert. Furthermore, existing VLA approaches lack a reliable mechanism for detecting task completion, leading to redundant actions that reduce overall task success rates. To address these challenges, we propose World-Env, an RL-based post-training framework that replaces physical interaction with a low-cost, world model-based virtual simulator. World-Env consists of two key components: (1) a video-based world simulator that generates temporally consistent future visual observations, and (2) a vision-language model (VLM)-guided instant reflector that provides continuous reward signals and predicts action termination. This simulated environment enables VLA models to safely explore and generalize beyond their initial imitation learning distribution. Our method achieves notable performance gains with as few as five expert demonstrations per task. Experiments on complex robotic manipulation tasks demonstrate that World-Env effectively overcomes the data inefficiency, safety constraints, and inefficient execution of conventional VLA models that rely on real-world interaction, offering a practical and scalable solution for post-training in resource-constrained settings. Our code is available at https://github.com/amap-cvlab/world-env.
Figure 1: Comparison of three VLA training paradigms: (a) Imitation learning suffers from poor generalization under data scarcity. (b) Prior RL-based post-training methods require real-world interaction, which is often infeasible due to non-resettable state transitions (e.g., object drop or collision). (c) Our proposed World-Env enables post-training via simulated rollouts using a world model, eliminating the need for physical interaction and supporting safe, efficient exploration even with minimal expert demonstrations.
@@ 1 Introduction
Vision-Language-Action (VLA) models have emerged as a central paradigm for autonomous agents, enabling end-to-end mapping from high-level language instructions to low-level motor commands by integrating vision, language, and control. These models have demonstrated considerable promise in robotic manipulation (Kim et al., 2024; Black et al., 2024), autonomous driving (Yurtsever et al., 2020), and navigation (Hong et al., 2021). Most existing approaches rely on supervised fine-tuning through imitation learning, building upon pre-trained vision-language models (Touvron et al., 2023a) to align semantic intent with physical execution via cross-modal representations.
However, imitation learning methods (Kim et al., 2025) are inherently constrained by the limited availability of high-quality demonstrations. In many real-world scenarios, collecting diverse and safe human demonstrations is prohibitively expensive and often infeasible due to safety concerns and environmental complexity. Furthermore, such methods generalize poorly to novel tasks or unseen objects, and their performance degrades under few-shot conditions.
To overcome these shortcomings, recent works (Tan et al., 2025; Lu et al., 2025) have turned to reinforcement learning (RL) (Rafailov et al., 2023) to enable agents to learn through interaction. Current RL strategies fall into two categories. The first involves real-world learning with human feedback, which captures realistic environmental dynamics but suffers from non-resettable interactions, high trial costs, and limited reproducibility, rendering it unsuitable for safety-critical applications. The second relies on simulator-based learning, which avoids physical risks but introduces other challenges, including substantial development effort, limited sim-to-real transfer, and difficulty adapting to new objects or dynamic scene changes, thereby restricting its practical applicability.
@@
In this work, we introduce World-Env, a world model-based reinforcement learning framework that improves policy generalization under data scarcity while respecting real-world safety constraints, as shown in Figure 1. World-Env consists of two components. The first is a video-based world simulator that functions as an interactive future-frame predictor, synthesizing action-conditioned image sequences that capture post-interaction object states and surrounding scene structure. The second is a VLM-guided instant reflector that functions as a semantics-aware reward module. It provides continuous reward signals by evaluating the semantic alignment between predicted visual frames and the input language instruction. This assessment supports policy optimization and enables real-time detection of task completion. Upon confirming successful execution (e.g., when the goal state is reached), the reflector immediately terminates the action sequence to prevent redundant or disruptive subsequent actions. The framework delivers three principal benefits: (1) efficient generalization from minimal expert demonstrations, (2) safe controllability through risk-free virtual exploration, and (3) language-aligned termination via VLM-driven reasoning.
@@
  • We propose World-Env, a world model-based framework that enables low-cost, safe reinforcement learning post-training for VLA policies under extreme data scarcity, eliminating the need for real-world interaction.
  • World-Env integrates a video-based world simulator and a VLM-guided instant reflector to jointly provide temporally consistent visual observations and continuous reward signals, forming a self-contained virtual environment that supports effective policy exploration.
@@ 2 Related Work
Vision-Language-Action Models.
@@
Leveraging advancements in pre-trained vision foundation models (Radford et al., 2021; Oquab et al., 2023; Dosovitskiy et al., 2020), large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023b; Bai et al., 2023), and vision-language models (VLMs) (Alayrac et al., 2022; Li et al., 2022; 2023; Dai et al., 2023; Liu et al., 2023c; b), Vision-Language-Action (VLA) frameworks (Chi et al., 2023; Team et al., 2024; Kim et al., 2024; Bu et al., 2025) have emerged as a powerful approach for embodied intelligence. These models bridge the gap between high-dimensional sensory inputs and physical-world actions by processing multimodal signals, such as camera feeds, sensor data, and natural language instructions, and translating them into actionable outputs for robotic systems. DiffusionPolicy (Chi et al., 2023) proposes a diffusion-based policy that generates robot actions through a conditional denoising diffusion process in the action space, iteratively refining actions based on visual observations. OpenVLA (Kim et al., 2024) integrates robotic actions into a language modeling framework by mapping action sequences to discrete tokens within a large language model. The recent work OpenVLA-OFT (Kim et al., 2025) further converts discrete action sequences into continuous representations, achieving improved inference efficiency and task performance.
Reinforcement Learning for VLA Systems.
@@
Recent advances in reinforcement learning (RL) (Schulman et al., 2017; Rafailov et al., 2023) have demonstrated considerable potential in enhancing decision-making capabilities of large language models (LLMs) (Guo et al., 2025; Lightman et al., 2023; Ouyang et al., 2022; Lee et al., 2023). This progress has spurred growing interest in applying RL to Vision-Language-Action (VLA) systems (Tan et al., 2025; Lu et al., 2025; Chandra et al., 2025; Jiang et al., 2025a), where adaptive behavior is essential. Unlike supervised fine-tuning (SFT), which replicates static demonstrations, RL enables agents to refine policies through interaction, optimizing actions to maximize task-oriented rewards. This paradigm supports autonomous exploration, reward-driven adaptation, and improved robustness to partial observability, allowing VLA models to generalize to unseen scenarios while reducing reliance on costly human demonstrations. However, existing RL-based VLA methods typically require real-world interaction, which is often infeasible in high-risk or non-resettable scenarios.
World Models.
@@
World models (Assran et al., 2025; Ball et al., 2025), which are learned dynamical simulators that approximate environmental dynamics, have become foundational for safe and sample-efficient agent training. Early works (Hafner et al., 2019; 2020; 2025) demonstrated the effectiveness of model-based reinforcement learning in virtual environments, enabling agents to plan via imagined trajectories without real-world interaction. Recent advances, such as TD-MPC2 (Hansen et al., 2024), have improved scalability and policy learning efficiency across multi-task and multi-domain settings. Similarly, PWM (Georgiev et al., 2025) leverages pre-trained world models and first-order optimization to handle high-dimensional action spaces in multi-task RL. However, these methods typically rely on on-policy data, limiting their generalization to specific environments and downstream tasks. Building on advances in diffusion-based video generation (Ho et al., 2020; Rombach et al., 2022; Blattmann et al., 2023; Yang et al., 2024; Wan et al., 2025; Xing et al., 2024), we propose a framework that trains a world model on offline demonstration data and keeps it fixed during policy learning to predict future visual observations for VLA models. This decouples world model training from policy exploration, enabling broader applicability in resource-constrained or high-risk scenarios.
@@ 3 Preliminary
Vision-Language-Action Models.
Vision-language-action (VLA) models bridge natural language instructions with robotic control by translating semantic goals into low-level actions while grounding language in multimodal observations. Following recent VLA frameworks such as OpenVLA-OFT (Kim et al., 2025), the policy is implemented as a deterministic mapping that leverages a pretrained vision-language model to extract multimodal features, followed by a lightweight action head for continuous control. Specifically, given a history of RGB observations $\mathbf{o}_{1:t}$, proprioceptive states $\mathbf{s}_{1:t}$ (e.g., joint angles or end-effector poses), and a language instruction $\mathbf{g}$, the policy predicts a deterministic action $\mathbf{a}_{t}\in\mathbb{R}^{D}$ as:
@@
Reinforcement Learning.
Reinforcement learning (RL) formulates decision-making as a Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space (comprising visual observations $\mathbf{o}_{t}$ and proprioceptive states $\mathbf{s}_{t}$), $\mathcal{A}$ is the action space (e.g., continuous control commands $\mathbf{a}_{t}\in\mathbb{R}^{D}$), $\mathcal{P}$ denotes the transition dynamics, $\mathcal{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. The objective is to learn a policy $\pi_{\theta}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$ that maximizes the expected return:
$J(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right],$
where $r_{t}=\mathcal{R}(\mathbf{o}_{1:t},\mathbf{g})$. In practice, policy gradient methods often introduce stochasticity during training to enable exploration. The policy is updated using gradients of the form:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\mathbf{a}_{t}\sim\pi_{\theta}}\left[\nabla_{\theta}\pi_{\theta}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})\cdot A(\mathbf{o}_{1:t},\mathbf{a}_{1:t})\right],$ (2)
@@ 4 Method
Figure 2: Overview of World-Env. Our framework comprises: (1) a Training Data Strategy that augments human demonstrations trajectories with VLA self-explored trajectories to train the World-Env Simulator; (2) an Optimization Loop where the VLA model generates actions, the simulator predicts future observations, and the World-Env Reflector generates feedback; and (3) Reward & End Signal provides trajectory-wise reward and end signals for RL optimization.
@@
Figure 2 presents the overview of our framework. Prior VLA approaches (Kim et al., 2024; 2025) typically rely on either real-world interaction or conventional simulators to provide observations for action prediction. In contrast, our framework eliminates the need for physical interaction by leveraging a video-based world simulator that generates temporally consistent future visual observations at low cost. Specifically, the deterministic VLA policy $\pi_{\theta}$ maps the current RGB observation $\mathbf{o}_{t}$, language instruction $\mathbf{g}$, and proprioceptive state $\mathbf{s}_{t}$ (comprising the 6D end-effector pose and 1D gripper state) to a continuous action $\mathbf{a}_{t}$. The next proprioceptive state $\mathbf{s}_{t+1}$ is then computed deterministically from $\mathbf{s}_{t}$ and $\mathbf{a}_{t}$ using forward kinematics. The world simulator takes the executed action $\mathbf{a}_{t}$ and the resulting proprioceptive state $\mathbf{s}_{t+1}$ as inputs and predicts the subsequent visual observation $\mathbf{o}_{t+1}$. This imagined observation, together with $\mathbf{s}_{t+1}$, is fed back into the VLA policy to predict the next action $\mathbf{a}_{t+1}$. The rollout terminates either when the maximum timestep is reached or when the VLM-guided instant reflector, which evaluates semantic alignment between the predicted visual trajectory and the language instruction, confirms task success and issues a termination signal. During training, we collect $N$ simulated trajectories from this virtual environment and use them for reinforcement learning (RL) optimization of the VLA policy within World-Env.
4.1 Video-Based World Simulator
Our world simulator is built upon the EVAC framework (Jiang et al., 2025b). During rollout, the simulator takes the executed action $\mathbf{a}_{t}$ and the resulting proprioceptive state $\mathbf{s}_{t+1}$ as inputs to predict the next visual observation $\mathbf{o}_{t+1}$. The proprioceptive state $\mathbf{s}_{t+1}$ comprises a 3D position vector $\mathbf{x}_{t+1}\in\mathbb{R}^{3}$, a 3D rotation vector $\mathbf{q}_{t+1}\in\mathbb{R}^{3}$ (represented in axis-angle format), and a 1D gripper state $p_{t+1}\in[0,1]$.
Following Jiang et al. (2025b), we render an action map by projecting the proprioceptive state $\mathbf{s}_{t+1}$ onto the image plane. This action map consists of a foreground marker indicating the projected pose and a black background to enhance visual contrast. The action map, together with the history observation, is injected into the EVAC world model as pixel-level conditioning. The EVAC model then generates the future observation $\mathbf{o}_{t+1}$ using a diffusion-based image generation module. For further architectural details, we refer readers to Jiang et al. (2025b).
@@ 4.1 Video-Based World Simulator
To train the world model, we find that relying solely on expert demonstrations from the LIBERO benchmark (Liu et al., 2023a) limits generalization to unseen state-action sequences. To address this, we augment the training data by enabling autonomous exploration in the LIBERO simulator. Specifically, we deploy the supervised fine-tuned OpenVLA-OFT policy (Kim et al., 2025) to predict actions and execute them in the simulator, which yields the corresponding next proprioceptive state $\mathbf{s}_{t+1}$ and observation $\mathbf{o}_{t+1}$. To further enhance data diversity, we introduce controlled stochasticity by training a scale head that predicts the log-scale parameter $\bm{\beta}_{t}$ of a Laplace distribution, with the OpenVLA-OFT action $\bm{\mu}_{t}$ as the location parameter: $\mathbf{a}_{t}\sim\text{Laplace}(\bm{\mu}_{t},\bm{\beta}_{t})$. These perturbed actions are executed to collect additional $(\mathbf{o}_{t},\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},\mathbf{o}_{t+1})$ transition tuples. Finally, we combine these autonomously collected trajectories with the original human-demonstrated successful trajectories from LIBERO (Liu et al., 2023a) to form a diverse and robust training dataset for the world simulator. Additional analysis of the world simulator and network architecture are provided in the supplementary material.

### v5

@@ Abstract
Vision-Language-Action (VLA) models trained via imitation learning suffer from significant performance degradation in data-scarce scenarios due to their reliance on large-scale demonstration datasets. Although reinforcement learning (RL)-based post-training has proven effective in addressing data scarcity, its application to VLA models is hindered by the non-resettable nature of real-world environments. This limitation is particularly critical in high-risk domains such as industrial automation, where interactions often induce state changes that are costly or infeasible to revert. Furthermore, existing VLA approaches lack a reliable mechanism for detecting task completion, leading to redundant actions that reduce overall task success rates. To address these challenges, we propose World-Env, an RL-based post-training framework that replaces physical interaction with a low-cost world model-based virtual simulator. World-Env consists of two key components: (1) a physically-consistent world simulator that generates temporally consistent future visual observations, and (2) a vision-language model (VLM)-guided instant reflector that provides continuous reward signals and predicts action termination. This simulated environment enables VLA models to safely explore and generalize beyond their initial imitation learning distribution. Our method achieves notable performance gains with as few as five expert demonstrations per task. Experiments on complex robotic manipulation tasks demonstrate that World-Env effectively overcomes the data inefficiency, safety constraints, and inefficient execution of conventional VLA models that rely on real-world interaction, offering a practical and scalable solution for post-training in resource-constrained settings. Our code is available at https://github.com/amap-cvlab/world-env.
Figure 1: VLA exploration in world model and performance comparison on LIBERO benchmarks.
[…]
@@ 1 Introduction
Vision-Language-Action (VLA) models have emerged as a central paradigm for autonomous agents, enabling end-to-end mapping from high-level language instructions to low-level motor commands by integrating vision, language, and control. These models have demonstrated considerable promise in robotic manipulation [27, 6], autonomous driving [56, 31, 55, 49, 57], and navigation [21, 58]. Most existing approaches rely on supervised fine-tuning through imitation learning, building upon pre-trained vision-language models [46] to align semantic intent with physical execution via cross-modal representations.
However, imitation learning methods [26] are inherently constrained by the limited availability of high-quality demonstrations. In many real-world scenarios, collecting diverse and safe human demonstrations is prohibitively expensive and often infeasible due to safety concerns and environmental complexity. Furthermore, such methods generalize poorly to novel tasks or unseen objects, and their performance degrades under few-shot conditions.
To overcome these shortcomings, recent works [44, 36] have turned to reinforcement learning (RL) [41] to enable agents to learn through interaction. Current RL strategies fall into two categories. The first involves real-world learning with human feedback, which captures realistic environmental dynamics but suffers from non-resettable interactions, high trial costs, and limited reproducibility, rendering it unsuitable for safety-critical applications. The second relies on simulator-based learning, which avoids physical risks but introduces other challenges, including substantial development effort, limited sim-to-real transfer, and difficulty adapting to new objects or dynamic scene changes, thereby restricting its practical applicability.
Figure 2: Comparison of three VLA training paradigms.
@@
In this work, we introduce World-Env, a world model-based reinforcement learning framework that improves policy generalization under data scarcity while respecting real-world safety constraints, as shown in Figure 2. World-Env consists of two components. The first is a physically-consistent world simulator that functions as an interactive future-frame predictor, synthesizing action-conditioned image sequences that capture post-interaction object states and surrounding scene structure. To ensure physical consistent of our world model, we propose a geometry-aware feature injection strategy, which leverages latent feature from VGGT [51] as additional condition. The second is a VLM-guided instant reflector that functions as a semantics-aware reward module. It provides continuous reward signals by evaluating the semantic alignment between predicted visual frames and the input language instruction. This assessment supports policy optimization and enables real-time detection of task completion. Upon confirming successful execution (e.g., when the goal state is reached), the reflector immediately terminates the action sequence to prevent redundant or disruptive subsequent actions.
@@
  • We propose World-Env, a framework that enables low-cost, safe reinforcement learning post-training for VLA policies under extreme data scarcity, eliminating the need for real-world interaction.
  • We propose a geometry-aware feature injection strategy via injecting latent feature from VGGT to ensure physical consistent of world model.
@@ 2 Related Work
Figure 3: Overview of World-Env. Our framework comprises: (1) a Training Data Strategy that augments human demonstrations trajectories with VLA self-explored trajectories to train the World Simulator; (2) an Optimization Loop where the VLA model generates actions, the simulator predicts future observations, and the Instant Reflector generates feedback; and (3) Reward & End Signal provides trajectory-wise reward and end signals for RL optimization.
@@
Vision-Language-Action Models. Leveraging advancements in pre-trained vision foundation models [40, 37, 14], large language models (LLMs) [8, 47, 4], and vision-language models (VLMs) [2, 30, 29, 13, 35, 34], Vision-Language-Action (VLA) frameworks [12, 45, 27, 9] have emerged as a powerful approach for embodied intelligence. Specifically, DiffusionPolicy [12] proposes a diffusion-based policy that generates robot actions through a conditional denoising diffusion process in the action space, iteratively refining actions based on visual observations. OpenVLA [27] integrates robotic actions into a language modeling framework by mapping action sequences to discrete tokens within a large language model. The recent work OpenVLA-OFT [26] further converts discrete action sequences into continuous representations, achieving improved efficiency and performance.
@@
Reinforcement Learning for VLA Systems. Recent advances in reinforcement learning (RL) [43, 41] have demonstrated considerable potential in enhancing decision-making capabilities of large language models (LLMs) [15, 32, 38, 28]. This progress has spurred growing interest in applying RL to Vision-Language-Action (VLA) systems [44, 36, 10, 24]. Unlike supervised fine-tuning (SFT), RL enables agents to refine policies through interaction. This paradigm supports autonomous exploration and improved robustness to partial observability, allowing VLA models to generalize to unseen scenarios while reducing reliance on costly human demonstrations. However, existing RL-based VLA methods typically require real-world interaction, which is often infeasible in high-risk scenarios.
@@
World Models. World models [3, 5], which are learned dynamical simulators that approximate environmental dynamics, have become foundational for safe and sample-efficient agent training. Early works [16, 17, 18] demonstrated the effectiveness of model-based reinforcement learning in virtual environments, enabling agents to plan via imagined trajectories without real-world interaction. However, these methods typically rely on on-policy data, limiting their generalization to specific environments and downstream tasks. Building on advances in diffusion-based video generation [20, 42, 7, 54, 50, 53], we propose a framework that trains a world model on offline demonstration data and keeps it fixed during policy learning to predict future visual observations for VLA models.
@@ 3 Preliminary
Vision-Language-Action Model. Vision-language-action (VLA) models bridge natural language instructions with robotic control by translating semantic goals into low-level actions while grounding language in multimodal observations. Following recent VLA frameworks such as OpenVLA-OFT [26], the policy is implemented as a deterministic mapping that leverages a pretrained vision-language model to extract multimodal features, followed by a lightweight action head for continuous control. Specifically, given a history of RGB observations $\mathbf{o}_{1:t}$, proprioceptive states $\mathbf{s}_{1:t}$ (e.g., joint angles or end-effector poses), and a language instruction $\mathbf{g}$, the policy predicts a action as:
@@
Reinforcement Learning. Reinforcement learning (RL) formulates decision-making as a Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state space (comprising visual observations $\mathbf{o}_{t}$ and proprioceptive states $\mathbf{s}_{t}$), $\mathcal{A}$ is the action space (e.g., continuous control commands $\mathbf{a}_{t}\in\mathbb{R}^{D}$), $\mathcal{P}$ denotes the transition dynamics, $\mathcal{R}$ is the reward function, and $\gamma\in[0,1]$ is the discount factor. The objective is to learn a policy $\pi_{\theta}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$ that maximizes the expected return:
$J(\pi_{\theta})=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right],$ (2)
where $r_{t}=\mathcal{R}(\mathbf{o}_{1:t},\mathbf{g})$. In practice, policy gradient methods often introduce stochasticity during training to enable exploration. The policy is updated using gradients:
$\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\mathbf{a}_{t}\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})\cdot A(\mathbf{o}_{1:t},\mathbf{a}_{1:t})\right],$ (3)
@@ 4 Method
Figure 4: Overview of physically-consistent world simulator.
@@
Figure 3 presents the overview of our framework. Prior VLA approaches [27, 26] typically rely on either real-world interaction or conventional simulators to provide observations for action prediction. In contrast, our framework eliminates the need for physical interaction by leveraging a physically-consistent world simulator that generates temporally consistent future visual observations at low cost. Specifically, the VLA policy $\pi_{\theta}$ maps the current RGB observation $\mathbf{o}_{t}$, language instruction $\mathbf{g}$, and proprioceptive state $\mathbf{s}_{t}$ (comprising the 6D end-effector pose and 1D gripper state) to a continuous action $\mathbf{a}_{t}$. The next proprioceptive state $\mathbf{s}_{t+1}$ is then computed deterministically using forward kinematics. The world simulator takes the resulting proprioceptive state $\mathbf{s}_{t+1}$ as input and predicts the subsequent visual observation $\mathbf{o}_{t+1}$. This imagined observation, together with $\mathbf{s}_{t+1}$, is fed back into the VLA policy to predict the next action $\mathbf{a}_{t+1}$. The rollout terminates either when the maximum timestep is reached or when the VLM-guided instant reflector, which evaluates semantic alignment between the predicted visual trajectory and the language instruction, confirms task success and issues a termination signal. During training, we collect $N$ simulated trajectories from this virtual environment and use them for RL optimization of the VLA policy within World-Env.
$\pi_{0}$ [6] 67.6 68.4 80.2 28.2 61.1
$\pi_{0}$+FAST [39] 59.2 76.8 59.2 24.8 55.0
OpenVLA [27] 73.2 55.0 82.4 32.2 60.7
UniVLA [9] 82.0 76.2 84.4 56.4 74.75
OpenVLA-OFT [26] 84.0 74.2 84.2 57.0 74.85
OpenVLA-OFT + Post-training (Ours) 86.4 86.6 87.6 57.8 79.6
Figure 5: Comparison with SFT on multi-goal tasks. Note that all results are collected every 5 training steps for three distinct goals.
4.1 Physically-Consistent World Simulator
Figure 4 presents an overview of our world simulator, which is designed to predict future observations conditioned on a given action sequence. Akin to [25], we first translate the predicted action into a proprioceptive state $\mathbf{s}_{t+1}$ via forward kinematics, and project it onto the image plane to generate an action map. This action map comprises a foreground marker that encodes the projected pose (e.g., position and orientation) and a uniformly black background, which serves to maximize visual contrast and minimize interference with the underlying scene content. The resulting action map is then combined with a historical observation sampled from a memory bank and jointly injected into a U-Net based denoising diffusion network as pixel-level conditioning signals. To ensure that the synthesized future frames remain physically plausible and geometrically coherent with the reference observation, we introduce a geometry-aware feature injection mechanism. Specifically, we extract complementary features from two pre-trained encoders: (i) VGGT [51], which excels at preserving fine-grained geometric structures and spatial layouts of the reference image; and (ii) CLIP [40], which captures high-level semantic and contextual information. These features are integrated into the denoising U-Net through cross-attention layers at multiple resolutions, enabling the model to simultaneously respect both local geometric fidelity and global semantic consistency during future frame synthesis. This dual-path injection strategy not only enhances the realism of generated observations but also improves temporal coherence and physical plausibility in long-horizon predictions.
@@ 4.1 Physically-Consistent World Simulator
To train the world model, we find that relying solely on expert demonstrations from the LIBERO benchmark [33] limits generalization to unseen state-action sequences. To address this, we augment the training data by enabling autonomous exploration in the LIBERO simulator. Specifically, we deploy the supervised fine-tuned OpenVLA-OFT policy [26] to predict actions and execute them in the simulator, which yields the corresponding next proprioceptive state $\mathbf{s}_{t+1}$ and observation $\mathbf{o}_{t+1}$. To further enhance data diversity, we introduce controlled stochasticity by training a scale head that predicts the log-scale parameter $\boldsymbol{\beta}_{t}$ of a Laplace distribution, with the OpenVLA-OFT action $\boldsymbol{\mu}_{t}$ as the location parameter: $\mathbf{a}_{t}\sim\text{Laplace}(\boldsymbol{\mu}_{t},\boldsymbol{\beta}_{t})$. These perturbed actions are executed to collect additional $(\mathbf{o}_{t},\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{s}_{t+1},\mathbf{o}_{t+1})$ transition tuples. Finally, we combine these autonomously collected trajectories with the original human-demonstrated successful trajectories from LIBERO [33] to form a diverse and robust training dataset for the world simulator. Additional analysis of the world simulator are provided in the supplementary material.